OpenStudy (anonymous):

True or False: (1/2)(2n + 2) + -2 = n

OpenStudy (mathstudent55):

Add 2 to both sides. What do you get?

OpenStudy (mathstudent55):

Then multiply both sides by 2.

OpenStudy (anonymous):

this question confuses me so much

OpenStudy (mathstudent55):

We're going to solve this as an equation. To get rid of the -2 on the left side, add 2 to both sides.

OpenStudy (mathstudent55):

(1/2)(2n + 2) + -2 = n +2 + 2 ---------------------------- (1/2)(2n + 2 = n + 2

OpenStudy (anonymous):

we have : (1/2)(2n + 2) + -2 = 1/2 * 2n + 1/2 * (-2)= n+1-2=n-1 Got it ?:)

OpenStudy (anonymous):

thanks you! so its false then?

OpenStudy (mathstudent55):

Since the left side is being multipled by 1/2, which is the same as dividing by 2, we're going to multiply both sides by 2: 2n + 2 = 2n + 4 Subtract 2n from both sides: 2 = 4 which is a false statement, so it's false.

OpenStudy (anonymous):

and this is true right? True or False: (-5 + n) + (5 + -n) = 0 because (-5 + 5) + (n + -n)= 0 correct?

OpenStudy (mathstudent55):

Yes, this is true. You are correct.